# Technical report on label-informed logit redistribution for better domain generalization in low-shot classification with foundation models

### v2

@@ Abstract
Confidence calibration is an emerging challenge in real-world decision systems based on foundations models when used for downstream vision classification tasks. Due to various reasons exposed, logit scores on the CLIP head remain large irrespective of whether the image-language pairs reconcile. It is difficult to address in data space, given the few-shot regime. We propose a penalty incorporated into loss objective that penalizes incorrect classifications whenever one is made during finetuning, by moving an amount of log-likelihood to the true class commensurate to the relative amplitudes of the two likelihoods. We refer to it as confidence misalignment penalty (CMP). Extensive experiments on $12$ vision datasets and $5$ domain generalization datasets supports the calibration performance of our method against stat-of-the-art. CMP outperforms the benchmarked prompt learning methods, demonstrating average improvement in Expected Calibration Error (ECE) by average $6.01$%, $4.01$ at minimum and $9.72$% at maximum.
Keywords Foundation models  $\cdot$ Prompt learning  $\cdot$ Confidence Calibration
@@ 1 Introduction
Recent developments in foundation models, such as CLIP Radford et al. (2021), ALIGN Jia et al. (2021), and Flamingo Alayrac et al. (2022), have demonstrated remarkable performance across a wide range of tasks by leveraging pre-trained multimodal representations, including few-shot and zero-shot classification with tunable language prompts. However, when these foundation models are used for downstream tasks (e.g., classification or retrieval tasks), the setup is observed to frequently lead to poor calibration and produces overconfident misclassified predictions Ye et al. (2023); Tu et al. (2024a) on the downstream task. This leads to problems such as an eventual reduction in generalization and the ascribing of undesirably high confidence to misclassified and out-of-distribution examples. This has in practice resulted in decisions with ethical, equity and health & safety consequences Firoozi et al. (2023), thereby eroding the trustworthiness of this iteration of AI for the general public. This could be viewed as a special case of misalignment, where the behavior of foundation models deviates from its objective, or expectation of users or society. The major cause for this is training on massive amount of data without any inherent understanding of task requirement or alignment to specific guideline or safety conditions Ouyang et al. (2022); Achiam et al. (2023).
Hypotheses
The probability mass that is put on a certain class choice could be interpreted as the relative confidence it has on that class explaining the example Guo et al. (2017); Li et al. (2024). We notice a few possible problems with this scheme:
@@
    Softmax logits may be viewed to be optimised for the classification problem in foundation models via a variational formulation (App. A) of the problem of estimating $P(\textbf{y}|\textbf{x};\theta)$. The high-bias from the variational estimate is antagonistic to the maximum entropy principle (MEP) (ref. exponential families) Jaynes (1957), which postulates that a more uniform, highest-entropy distribution best explains observed data. This leads to the neural network - softmax classifier pair make overconfident misclassifications.
@@
    The optimization dynamics of contrastive learning lead to overconfidence. The objective function pushes the matched pair (image and text) embeddings into closer clusters together while separating the mismatched pairs. The alignment may result in highly concentrated clusters, causing similarity score for certain predictions to dominates the others. There is a parallel understanding of the CLIP loss undergoing neural collapse in the final phase of training Papyan et al. (2020). In that regime, it is noticed that intra-class examples close in or collapse towards the class mean, this speed moderated by the temperature of the softmax. The rescaling by the softmax to $1$ may further reduces variability. The higher and narrower peak encourage over-confidence, regardless of the correctness of the classification. Small amounts of embedding misalignment leads to a large confidence perturbation for the downstream task Desai and Durrett (2020); Rudner et al. (2024).
  3. 3.
    The rigidity of representations in frozen foundational models could contribute to poor adaptation in downstream tasks because the embeddings remain quasi-static. They have to be exposed to larger amounts of finetuning data than afforded by few-shot finetuning to be optimized for task-specific nuances, limiting their ability to capture domain-specific features. While this rigidity may not directly cause overfitting in the traditional sense, but it can exacerbate overconfidence in incorrect predictions due to the model reliance on fixed, potentially suboptimal representations Bommasani et al. (2021); Geng et al. (2024).
Model uncertainty often arises under distribution shift/ domain shift, such as due to bad weather condition or poor visibility, then failing gracefully is a more autonomous option than to transferring control to a fallback system (i.e human) Filos et al. (2020). This work tunes the logits to be less wrong when misclassifying.
Contributions
This work makes the following major contributions:
  1. 1.
    We introduce a loss component that suppresses CLIP’s confidence on misclassifications by moderating the flow of posterior likelihood towards any posterior peak higher than that of the true class.
  2. 2.
    We use knowledge about incorrect predictions existing as labels, which is inherently invariant to transformations such as data augmentation or batch/layer-normalization.
  3. 3.
    We offer improvements in the predictive envelope of a network along the axes of network calibration, trustworthiness and generalisation, using metrics that provide direct measures of calibration; and derived metrics that provide a relative score to learner trustworthiness.
@@ 2 Related work
CLIP foundation model Radford et al. (2021), consist of an image encoder $f(x)$ and a text encoder $g(\textbf{t})$. Both encoders are trained using contrastive loss. The pre-trained CLIP provide support of zero-shot learning which can then be used for downstream tasks like prompt learning using fixed or hand-crafted prompts. Context Optimization (CoOp) Zhou et al. (2022a) introduced static learnable task specific prompts by replacing the manually designed prompts. CoOp provides two different prompt implementation i.e. Unified context vectors where prompts shared across classes and Class-specific context vectors provides individual prompts per class. But admittedly, CoOp struggles with generalization to unseen classes and distribution shift. Conditional Context Optimization (CoCoOp) Zhou et al. (2022b) generates dynamic prompts using light neural network adaptive to input image. The generated prompt are instance-specific dynamic prompts capable to address overfitting. Prompt distribution learning (ProDA) Lu et al. (2022) models a distribution of diverse prompts instead of individual prompts to generalise well to downstream tasks. ProDA learns a Gaussian distribution of classifiers weights drawn from diverse prompts. Diversity is introduced by varying class-token position in prompt.
@@
Confidence calibration in deep learning models. Extensive research focuses on confidence calibration methods to ensure model accuracy is aligned with its predicted confidence scores. These methods includes regularization based approaches, such as implicit regularization Ross and Dollár (2017) Baloch et al. (2019), $L_{2}$ regularization Guo et al. (2017) , and entropy regularization Pereyra et al. (2017) to align the model predicted accuracy with confidence. Some augmentation-based approaches, such as label-smoothing Müller et al. (2019) and mix-up training methods Thulasidasan et al. (2019); Zhang et al. (2022) are also explored for confidence calibration in deep learning models. Ovadia et al. (2019) empirically investigated the calibration of deep learning models under distribution shift. A comprehensive survey Wang (2023), which addresses the recent development about calibration for deep learning models.
@@
Confidence calibration in foundation models. Pandey et al. (2024) addresses the issue of under-confidence in pre-trained foundation models when these models are fine-tuned for downstream tasks in few-shot settings. Similarly, Murugesan et al. (2025) highlights the miscalibration problem in CLIP-based adaptive approaches. Their proposed method empirically demonstrates that CLIP-based adaptive approaches, such as prompt learning, adapters, parameter-efficient fine-tuning (PEFT), and test-time adaptation, are prone to miscalibration in zero-shot settings under distribution shifts. Furthermore, Tu et al. (2024b) investigates the prediction uncertainty of the foundation model CLIP for image classification tasks, focusing on variations in visual factors.
@@
In this work, we propose a regularization-based approach to ensure that the predictive probabilities of foundation models align with confidence scores in a prompt-learning setting under distribution shifts.
@@ 3 Method
A fundamental issue with the otherwise highly effective pairing of softmax-based classifiers and cross-entropy loss is prediction overconfidence and underconfidence, where the softmax logits, which are unnormalized log probabilities at the output, tend to overestimate or underestomate the true classification accuracy. Foundation models such as CLIP use contrastive loss to align text and image embeddings Radford et al. (2021).
The CLIP loss $\mathcal{L}_{\text{c}}$ can be defined as:
$\mathcal{L}_{\text{c}}=\frac{1}{2}\left(\mathcal{L}_{\text{txt}}+\mathcal{L}_{\text{img}}\right),$ (1)
while,
$\mathcal{L}_{\text{txt}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\left(\text{sim}(\mathbf{t}_{i},\mathbf{i}_{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(\mathbf{t}_{i},\mathbf{i}_{j})/\tau\right)},$ (2)
$\mathcal{L}_{\text{img}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\left(\text{sim}(\mathbf{i}_{i},\mathbf{t}_{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(\mathbf{i}_{i},\mathbf{t}_{j})/\tau\right)}$ (3)
@@
CLIP’s contrastive loss for aligning image-text pairs embeddings which introduces bias to posterior probability $P(\mathbf{t}\mid\mathbf{i};\theta)$ and $P(\mathbf{i}\mid\mathbf{t};\theta)$ towards low entropy distribution, conflicting with the Maximum Entropy Principle (MEP).
@@
where,
$P(\mathbf{t}|\mathbf{i};\theta)=\frac{\exp(\text{sim}(\mathbf{i},\mathbf{t})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(\mathbf{i},\mathbf{t}_{j})/\tau)}.$
where $\text{sim}(\mathbf{i},\mathbf{t})$ is pair cosine similarity, $\tau$ is learned temperature parameter and N is batch size.
Posterior distribution ’s $P(\mathbf{t}\mid\mathbf{i};\theta)$ entropy is given by:
@@
CLIP contrastive loss minimizes $\mathcal{L}_{c}$ by maximizng the cosine similarity for all correct pairs while minimizing it for all negative pairs. It pushes $P(\mathbf{t}_{i}|\mathbf{i})l\rightarrow 1$ and $P(\mathbf{t}_{j}|\mathbf{i})\rightarrow 0$ for $j\neq i$. As per def. of MEP Jaynes (1957), more uniform, highest-entropy distribution best explains observed data.
@@
MEP normalization constraint ($\sum_{j}P(\mathbf{t}_{j}|\mathbf{i})=1$) favors more uniform probability distribution where all classes shares non-negligible probabilities under uncertainty.
CLIP loss $\mathcal{L}_{c}$ penalize the similarity score $\text{sim}(\mathbf{i},\mathbf{t}_{j})$ that is not the highest. As model minimizes $\mathcal{L}_{\text{c}}$, the probabilities $P(\mathbf{t}_{j}|\mathbf{i})$ for $j\neq i$ approach zero:
$P(\mathbf{t}_{j}|\mathbf{i})\propto\exp(\text{sim}(\mathbf{i},\mathbf{t}_{j})/\tau),\quad\text{where }\text{sim}(\mathbf{i},\mathbf{t}_{j})\ll\text{sim}(\mathbf{i},\mathbf{t}_{i}).$
which results in:
$H(P(\mathbf{t}|\mathbf{i};\theta))\rightarrow 0.$
Thus probability distributions become highly peaked which contradict MEP priniciple. ∎
@@
For well-calibrated prediction,the sofmax probability $P(x,y^{\prime})$ assigned to the correct predicted class $y^{\prime}$ must be higher than the incorrect classes $(P(x,y_{j}),j\neq y^{\prime})$. Therefore, we re-introduce a fraction of likelihood to penalize the contrastive loss, the confidence misalignment penalty (CMP).
$\text{CMP}=\frac{P(x,y^{\prime})}{\sum_{y_{j}\neq y^{\prime}\,:\,P(x,y_{j})>P(x,y^{\prime})}P(x,y_{j})},$ (4)
where $P(x,y^{\prime})=\frac{\exp(s(x,y^{\prime}))}{\sum_{j}\exp(s(x,y_{j}))}$ is softmax probability for correct pair and $P(x,y_{j})=\frac{\exp(s(x,y_{j}))}{\sum_{k}\exp(s(x,y_{k}))}$ represents the softmax probabilities for all incorrect pairs $(x,y_{j})$ where $y_{j}\neq y^{\prime}$.
@@
$\text{CMP}\leq 1$ thereby bounding the confidence by redistributing the probability mass from overconfident incorrect predictions to the correct prediction. (see proof detail A.1)
@@
$\mathcal{L}_{\text{CMP}}=\frac{1}{N}\sum_{i=1}^{N}\frac{P(x_{i},y_{i})}{\sum_{y_{j}\neq y_{i},\,P(x_{i},y_{j})>P(x_{i},y_{i})}P(x_{i},y_{j})+\epsilon}.$ (5)
Proposition 3.2.
The final combined loss $\mathcal{{L}_{\text{final}}}$ has a well-defined lower bound and ensures convergence during optimization. (see proof detail A.2)
@@
The CMP penalty for correctly classified pairs is $\mathcal{L}_{\text{CMP}}(x,y^{\prime})\approx 0$, while for misclassified pairs, $\mathcal{L}_{\text{CMP}}(x,y^{\prime})>0$. (see proof detail A.3)
@@
The gradient of the CMP loss $\mathcal{L}_{\text{CMP}}$ with respect to logits is bounded and ensures stable optimization. (see proof detail A.4)
The introduced CMP addresses the critical challenge of miscalibration in foundation models, by redistributing the probability mass from incorrect prediction to correct one. CMP integration into CLIP ensures reliable predictions which are aligned with true probability. Furthermore, theoretical guarantees like confidence bound, gradient stability, convergence and selective penalty makes the propsed method robust for use in real-world Vigneron et al. (2010) applications where trust is in predictions is critical.
Illustrative example
We illustrate confidence bleed to competing classes using one class 00 (Plane) in CIFAR10101010 as a toy example. The top left quadrant of Fig. 1 shows confidence by definition of logits, where classes besides class 00 have reasonably high bars, most notably class $8$ (Ship) and others. On its right is the equivalent logit vector for our method. This result carries over upon aggregation over the entire testing set, as shown in the respective quadrants on the bottom row, where correctly classified examples (in blue) versus misclassifications (in yellow). Our penalty introduces distinct variation between the highly-confident correct predictions and the low-confidence wrong predictions.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 1: Confidence calibration of CLIP in zero-shot setting for CIFAR10101010 dataset. The top row shows the confidence measurement for class 00 with contrastive loss (left) and confidence misalignment penalty loss (right). The bottom row denotes the confidence level as histogram for correct and incorrect predictions over test set where misalignment penalty shows improvement in calibration by showing logits being less confused.

### v3

# technical report on logit redistribution for better domain generalization in low-shot classification with foundation models

@@ Abstract
Confidence calibration is an emerging challenge in real-world decision systems that repurpose foundations models for downstream vision classification tasks. Due to various reasons such as inherent biases in contrastive pre-training, frozen embeddings and neural collapse in CLIP when post-training, logit scores on the CLIP head remain large irrespective of whether the image-language pairs reconcile. Ideally, they should be proportional to the amount of that reconciliation. This paper adaptively regulates that reconciliation. We propose a penalty incorporated into the oss objective that penalizes incorrect classifications whenever one is made during adaptation by moving an amount of log-likelihood to the true class commensurate to the relative amplitudes of the two likelihoods. We refer to the latter as the confidence misalignment penalty (CMP). Extensive experiments on $12$ vision datasets and $5$ domain generalization datasets supports the calibration performance of our method against stat-of-the-art. CMP outperforms the benchmarked prompt learning methods, demonstrating average improvement in Expected Calibration Error (ECE) by average 6.016.01%, $4.01$ % at minimum and 9.729.72% at maximum.
@@ 1 Introduction
Foundation models such as CLIP Radford et al. (2021), ALIGN Jia et al. (2021), and Flamingo Alayrac et al. (2022), have demonstrated remarkable performance across a wide range of tasks by leveraging pre-trained multimodal representations, including few-shot and zero-shot classification with tunable language prompts. However, when these foundation models are used for downstream tasks (e.g., classification or retrieval tasks), the setup is observed to frequently lead to poor calibration and produces overconfident misclassified predictions Ye et al. (2023); Tu et al. (2024a) on the downstream task. This leads to problems such as an eventual reduction in generalization and the ascribing of undesirably high confidence to misclassified and out-of-distribution examples. This has in practice resulted in decisions with ethical, equity and health & safety consequences Firoozi et al. (2023), thereby eroding the trustworthiness of this iteration of AI for the general public. This could be viewed as a special case of misalignment, where the behavior of foundation models deviates from its objective, or expectation of users or society. The major cause for this is post-training on data without an alignment of the downstream task to the often-undisclosed distribution of data in pre-training or to specific guideline or safety conditions Ouyang et al. (2022); Achiam et al. (2023). We understand the problem to have emerged from a mix of factors: 1) a discounting of the maximum entropy theorem (Sec. 3 Theorem 3.1), 2) neural collapse at post-training time (Sec. 3), and 3) upstream non-plasticity embedding under low-shot exposure to downstream data. We further specify the problem of post-training under the reality of distributional divergence. Model uncertainty often arises under distribution or domain shift, such as due to bad weather condition or poor visibility, then failing gracefully is a more autonomous option than to transferring control to a fallback system (i.e human) Filos et al. (2020). This work tunes the logits to be less wrong when misclassifying. In this contribution, we gather the amount of confidence put on the modal CLIP logit when a misclassification has occurred, reassign the difference to the correct text-image pair, therefore explicitly forcing parameter updates to fix the most broken part of the prediction distribution for a given example.
The work subscribes within the body of work on model uncertainty that is championed by Guo et al. (2017)’s model calibration with a focus on the empirical distribution of the predictions $\hat{y}$. It explores the problem from first principles of modeling the output distribution, and fixes a central problem that increases a more realist estimate of confidence, and increases system trust.
Contributions
@@
    We propose a loss component that suppresses CLIP’s confidence on mismatched language–image pairings by moderating the posterior flow whenever an incorrect class attains a higher peak than the true class.
@@
    We show through extensive experiments that CMP consistently improves both accuracy and calibration across few-shot and distribution-shift benchmarks.
Illustrative example
We illustrate undue confidence attribution to competing classes using one class 0 (Plane) in CIFAR1010 as a toy example. The top left quadrant of Fig. 1 shows confidence by definition of logits, where classes besides class 0 have reasonably high bars, most notably class $8$ (Ship) and others. On its right is the equivalent logit vector for our method. This result carries over upon aggregation over the entire testing set, as shown in the respective quadrants on the bottom row, where correctly classified examples (in blue) versus misclassifications (in yellow). Our penalty introduces distinct variation between the highly-confident correct predictions and the low-confidence wrong predictions.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 1: Confidence calibration of CLIP in low-shot (4-shot) settings for CIFAR1010 dataset produced using the CoOp pipeline. The top row shows the confidence measurement for class 0 with contrastive loss (left) and confidence misalignment penalty loss (right). The bottom row denotes the confidence level as histogram for correct and incorrect predictions over test set where misalignment penalty shows improvement in calibration by showing logits being less confused.
@@ 2 Related work
The CLIP foundation model Radford et al. (2021), consist of an image encoder $f(x)$ and a text encoder $g(\textbf{t})$. Both encoders are jointly trained using contrastive loss. The pre-trained CLIP provide support of zero-shot learning which can then be used for downstream tasks like prompt learning using fixed or hand-crafted prompts. Context Optimization (CoOp) Zhou et al. (2022a) introduced learnable task-specific prompts by replacing the manually designed prompts. CoOp provides two different prompt implementation i.e. Unified context vectors where prompts shared across classes and Class-specific context vectors provides individual prompts per class. But admittedly, CoOp struggles with generalization to unseen classes and distribution shift. Conditional Context Optimization (CoCoOp) Zhou et al. (2022b) generates dynamic prompts using light neural network adaptive to input image. The generated prompt are instance-specific dynamic prompts capable to address overfitting. Prompt distribution learning (ProDA) Lu et al. (2022) models a distribution of diverse prompts instead of individual prompts to generalise well to downstream tasks. ProDA learns a Gaussian distribution of classifiers weights drawn from diverse prompts. Diversity is introduced by varying class-token position in prompt.
@@
Confidence calibration in deep learning models. Extensive research focuses on confidence calibration methods to ensure model accuracy is aligned with its predicted confidence scores. These methods includes regularization based approaches, such as implicit regularization Ross and Dollár (2017) Baloch et al. (2019), $L_{2}$ regularization Guo et al. (2017) , and entropy regularization Pereyra et al. (2017) to align the model predicted accuracy with confidence. Some augmentation-based approaches, such as label-smoothing Müller et al. (2019) and mix-up training methods Thulasidasan et al. (2019); Zhang et al. (2022) are also explored for confidence calibration in deep learning models. Mukhoti et al. (2020) use focal loss to uniformly down-weight the confidence prediction for all samples while $CMP$ which penalize misclassified only. Liu et al. (2022) applies margin smoothing by assuming uniform label noise and requires tuning margin hyperparameters. $CMP$ automatically adapts penalty strength via relative logit amplitudes, requiring no margin tuning.
@@
Confidence calibration in foundation models. Pandey et al. (2024) addresses the issue of under-confidence in pre-trained foundation models when these models are fine-tuned for downstream tasks in few-shot settings. Similarly, Murugesan et al. (2025) highlights the miscalibration problem in CLIP-based adaptive approaches. Their proposed method empirically demonstrates that CLIP-based adaptive approaches, such as prompt learning, adapters, parameter-efficient fine-tuning (PEFT), and test-time adaptation, are prone to miscalibration in zero-shot/few-shot settings under distribution shifts. Furthermore, Tu et al. (2024b) investigates the prediction uncertainty of the foundation model CLIP for image classification tasks, focusing on variations in visual factors.
@@
In this work, we propose a regularization-based approach CMP which focus on low-shot regimes where validation set is infeasible to ensure that the predictive probabilities of foundation models align with confidence scores in a prompt-learning setting under distribution shifts.
@@ 3 Method
A fundamental issue with the otherwise highly effective pairing of softmax-based classifiers and cross-entropy loss is prediction overconfidence and underconfidence, where the softmax logits, which are unnormalized log probabilities at the output, tend to overestimate or underestomate the true classification accuracy. Foundation models such as CLIP use contrastive loss to align text and image embeddings Radford et al. (2021). The CLIP loss $\mathcal{L}_{\text{c}}$ can be defined as: $\mathcal{L}_{\text{c}}=\frac{1}{2}\left(\mathcal{L}_{\text{txt}}+\mathcal{L}_{\text{img}}\right),$ $\mathcal{L}_{\text{txt}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\left(\text{sim}(\mathbf{t}_{i},\mathbf{i}_{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(\mathbf{t}_{i},\mathbf{i}_{j})/\tau\right)},$ $\mathcal{L}_{\text{img}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\left(\text{sim}(\mathbf{i}_{i},\mathbf{t}_{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(\mathbf{i}_{i},\mathbf{t}_{j})/\tau\right)}$
@@
CLIP’s contrastive loss for aligning image-text pairs introduces bias to the posterior probabilities $P(\mathbf{t}\mid\mathbf{i};\theta)$ and $P(\mathbf{i}\mid\mathbf{t};\theta)$, pushing them towards low-entropy distributions, thereby conflicting with the Maximum Entropy Principle (MEP). This is reflected in the sharpness metric $\mathcal{S}$, defined as:
$\mathcal{S}=\frac{\max_{j}P(\mathbf{t}_{j}|\mathbf{i})}{\mathbb{E}_{\mathbf{t}_{j}\sim\mathcal{U}}[P(\mathbf{t}_{j}|\mathbf{i})]}=N\cdot P(\mathbf{t}_{i}|\mathbf{i})$
where CLIP optimization drives $\mathcal{S}\rightarrow N$ (fully peaked), while MEP favors $\mathcal{S}=1$ (uniform). Furthermore, the entropy $H(P(\mathbf{t}|\mathbf{i};\theta))$ is bounded by:
$H(P(\mathbf{t}|\mathbf{i};\theta))\leq\log N-\frac{N}{N-1}\exp\left(-\frac{2\Delta_{\text{sim}}}{\tau}\right),$
with $\Delta_{\text{sim}}=\text{sim}(\mathbf{i},\mathbf{t}_{i})-\max_{j\neq i}\text{sim}(\mathbf{i},\mathbf{t}_{j})$ is the similarity gap. As the training progresses and $\Delta_{\text{sim}}\to\infty$, the bound forces $H(P)\to 0$, directly opposing MEP’s principle.
@@
where: $P(\mathbf{t}|\mathbf{i};\theta)=\frac{\exp(\text{sim}(\mathbf{i},\mathbf{t})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(\mathbf{i},\mathbf{t}_{j})/\tau)}.$ Here, $\text{sim}(\mathbf{i},\mathbf{t})$ is the cosine similarity, $\tau$ is a learned temperature parameter, and $N$ is the batch size.
The entropy of $P(\mathbf{t}|\mathbf{i};\theta)$ is:
@@
CLIP’s loss minimization forces $P(\mathbf{t}_{i}|\mathbf{i})\rightarrow 1$ (correct pair) and $P(\mathbf{t}_{j}|\mathbf{i})\rightarrow 0$ (negative pairs, $j\neq i$), reducing entropy $H(P)\rightarrow 0$. This behavior is quantified by entropy bound which is derived from dominance of correct pair similarity:
$P(\mathbf{t}_{j}|\mathbf{i})\leq\frac{\exp(-\Delta_{\text{sim}}/\tau)}{1+(N-1)\exp(-\Delta_{\text{sim}}/\tau)},\quad j\neq i$
leading to the upper bound on entropy.
According to the Maximum Entropy Principle (MEP) Jaynes (1957), the least biased distribution that best represents uncertain data is the one with maximum entropy:
@@
which, under normalization ($\sum_{j}P(\mathbf{t}_{j}|\mathbf{i})=1$), favors a uniform distribution where all outcomes have non-negligible probability.
To quantify the deviation from MEP, we adapt sharpness metric $\mathcal{S}$ inspired by distributional peakedness measures Theodoridis (2015); Guo et al. (2017); Tian (2022):
$\mathcal{S}=\frac{\max_{j}P(\mathbf{t}_{j}|\mathbf{i})}{\mathbb{E}_{\mathbf{t}_{j}\sim\mathcal{U}}[P(\mathbf{t}_{j}|\mathbf{i})]}=N\cdot P(\mathbf{t}_{i}|\mathbf{i})$
since $\mathbb{E}[P(\mathbf{t}_{j}|\mathbf{i})]=\frac{1}{N}$ (due to uniform expectation under no assumptions).
Under CLIP optimization:
@@
whereas MEP requires $\mathcal{S}=1$ (uniform distribution).
Thus, CLIP’s loss minimization directly opposes MEP through both metrics:
$\mathcal{S}\rightarrow N\quad\text{and}\quad H(P)\rightarrow 0.$
We emphasize that MEP is applied selectively. CLIP’s contrastive loss drives all predictions toward peaked distributions ($S\rightarrow N$), but this is only problematic for misclassified examples where high confidence is unjustified. For correct predictions, peaked distributions (S → N) are desirable. CMP therefore applies MEP-inspired regularization only when $P(x,y^{\prime})>P(x,y)P(x,y^{\prime})$ (misclassification), preserving appropriate confidence for correct predictions while reducing overconfidence for errors.
However, these models are prone to confidence not being proportionate to predictions Wen et al. . The important remark is that for a wrong prediction $\hat{y}$, any class that is not the true image class $y^{\prime}$“steals” probability mass from $y^{\prime}$. Indeed, even for a correct prediction, this mass is spread to possibly all competing class choices. This amount of lost likelihood gives us a measure of confidence lost in the prediction. We call the quantity, confidence misalignment (CM).
For well-calibrated prediction,the sofmax probability $P(x,y)$ assigned to the correct predicted class $y$ must be higher than the incorrect classes $(P(x,y^{\prime}),y\neq y^{\prime})$. Therefore, we re-introduce a fraction of likelihood to penalize the contrastive loss, the confidence misalignment penalty (CMP)  [^1] We fondly like to refer to the model’s reflexive epistemics and do a wordplay with the name of the penalty. Referring to to the social psychologists Dunning and Kruger’s work Kruger and Dunning (1999), we informally call our loss the Dunning-Kruger Loss (the DKL), reading it as the reflected anagram (i.e. a guided redistribution) of KLD, the Kullback-Leibler Divergence, which fixates on a high-bias region around the mode..
$\text{CMP}=\frac{P(x,y)}{\sum_{y^{\prime}\neq y\,:\,P(x,y^{\prime})>P(x,y)}P(x,y^{\prime})},$ (1)
where $P(x,y)=\frac{\exp(s(x,y))}{\sum_{j}\exp(s(x,y))}$ is softmax probability for correct pair and $P(x,y^{\prime})=\frac{\exp(s(x,y^{\prime}))}{\sum_{k}\exp(s(x,y_{k}))}$ represents the softmax probabilities for all incorrect pairs $(x,y^{\prime})$ where $y\neq y^{\prime}$.
@@
$\text{CMP}\leq 1$ thereby bounding the confidence by redistributing the probability mass from overconfident incorrect predictions to the correct prediction. (see proof detail B.1)
@@
$\mathcal{L}_{\text{CMP}}=\frac{1}{N}\sum_{i=1}^{N}\frac{P(x_{i},y_{i})}{\sum_{y^{\prime}\neq y_{i},\,P(x_{i},y^{\prime})>P(x_{i},y_{i})}P(x_{i},y^{\prime})+\epsilon}.$ (2)
@@
The CMP penalty for correctly classified pairs is $\mathcal{L}_{\text{CMP}}(x,y)\approx 0$, while for misclassified pairs, $\mathcal{L}_{\text{CMP}}(x,y^{\prime})>0$. (see proof detail B.2)
@@
The gradient of the CMP loss $\mathcal{L}_{\text{CMP}}$ with respect to logits is bounded and ensures stable optimization. (see proof detail B.3)
The introduced CMP addresses the critical challenge of miscalibration in foundation models, by redistributing the probability mass from incorrect prediction to correct one. CMP integration into CLIP ensures reliable predictions which are aligned with true probability. Furthermore, theoretical guarantees like confidence bound, gradient stability, convergence and selective penalty makes the proposed method robust for use in real-world applications where trust is in predictions is critical.